A motorcycle breaks down and the rider must walk the rest of the way to work. The motorcycle was being driven at 45mph, and the rider walks at a speed of 6 mph. The distance from home to work is 25 miles, and the total for the trip was 2 hours. How far did the motorcycle go before it broke down?
Answer by stanbon(48558)

A motorcycle breaks down and the rider must walk the rest of the way to work. The motorcycle was being driven at 45mph, and the rider walks at a speed of 6 mph. The distance from home to work is 25 miles, and the total for the trip was 2 hours. How far did the motorcycle go before it broke down?
----------------------
Motorcycle DATA:
rate = 45 mph ; distance = x miles; time = x/45 hrs.
---------------------
Walking DATA:
rate = 6 mph ; distance = 25-x miles ; time = (25-x)/6 hrs.
----------------------
EQUATION:
time + time = 2 hrs
x/45 + (25-x)/6 = 2
6x + 45(25-x) = 2*6*45
6x + 25*45 - 45x = 12*45
-39x = -13*45
x = 15 miles (distance the motorcycle went)
==============================
Cheers,
Stan H.